# The Landscape of GPU-Centric Communication

### v2

@@ 3.2.4. GPUDirect Async
While previous GPUDirect technologies focused on improving the data path, GPUDirect Async optimizes the control path between the GPU and the NIC. Introduced in CUDA 8.0, it enables GPUs to initiate and synchronize network transfers, thereby reducing the CPU’s involvement in the critical path. GPUDirect Async works by having the CPU pre-register messages, which the GPU kernel can then trigger by ringing a doorbell on the NIC. As a result, the GPU can continue executing while the communication is being triggered, rather than needing to stop for the CPU to initiate the communication, as was previously necessary (Agostini et al., 2017, 2018).
@@ 3.3. GPUNetIO
GPUNetIO (Brosky, 2023a) is a technology solution proposed by NVIDIA as part of DOCA (Datacenter-On-a-Chip Architecture) (Corporation, 2023). DOCA is a full-stack software framework designed to facilitate the development of applications for NVIDIA BlueField Data Processing Units (DPUs) (Burstein, 2021). On non-RDMA networks GPUNetIO allows the GPU to send, receive, and process network packets (Brosky, 2023b, a). On RDMA networks (both RoCE and InfiniBand) (Corporation, 2023), from DOCA v2.7, GPUNetIO allows the GPU to execute RDMA send and receive not only on kernel boundaries, but at any point during the kernel execution. In a nutshell, GPUNetIO allows the GPU to interact with the NIC without any CPU intervention on the critical path.
@@ 3.4. Modern GPU-centric Interconnects
NVLink is a proprietary interconnect technology that facilitates high bandwidth and low latency direct access between NVIDIA GPUs. Its design addresses the bandwidth limitations of PCIe, which has been observed to be a transfer bottleneck in GPU-accelerated applications (Li et al., 2020; NVIDIA, 2017b). Table 3 presents the specifications of each generation of NVLink (Foley and Danskin, 2017; NVIDIA, 2023h; Li et al., 2020).
@@
Additionally, NVLink was also used to connect GPUs with the CPU for IBM Power8 and Power9 CPUs but with the introduction to Grace Hopper Superchip, NVLink is used as a Chip-to-Chip (C2C) interconnect with 900 GB/sec Bi-directional bandwidth (Li et al., 2020; NVIDIA, 2024b). Later with the introduction of Grace Blackwell Superchip, NVLink-C2C is used for connecting Grace CPU with 2 Blackwell GPUs with a total of 3.6 TB/sec bidirectional Bandwidth(NVIDIA, 2024a). Fifth-generation NVLink on NVIDIA Blackwell delivers 1.8TB/s bidirectional throughput per GPU, providing high-speed communication among up to 576 GPUs.
While direct GPU-to-GPU P2P communication had already been established over PCIe with GPUDirect P2P, it was heavily bottlenecked by the low bandwidths of PCIe. The introduction of NVLink optimized the bandwidth between NVIDIA GPUs, turning P2P communication into a viable mechanism for intra-node communication and shifting the data path heavily in favor of GPUs. A disadvantage of NVLink is that it is not self-routed meaning that if any two given GPUs do not have a direct NVLink connection communication will have to be routed through an intermediate GPU (Li et al., 2020). This limitation is overcome by NVSwitch (NVIDIA, 2023h), a backboard technology that can implement all-to-all connections between all GPUs. As an example, a DGX-2 node consists of 16 V100 GPUs that are all-to-all connected through NVLink and NVSwitch (NVIDIA, 2023c). Starting from the third generation, NVSwitch supports SHARP (NVIDIA, 2024c), which offloads allreduce operations to NVSwitch, allowing allreduce to operate at full line rate (Ishii and Wells, 2023).
AMD also offers a proprietary interconnect technology called xGMI/Infinity Fabric. However, it currently lacks a technology comparable to NVSwitch. While up to 8 MI300X GPUs can be connected in a full all-to-all mesh, NVSwitch can connect up to 64 GPUs. Recently, the Ultra Accelerator Link (UALink) consortium was established to develop a more open shared memory accelerator interconnect, compatible with multiple technologies and vendors (Morgan, 2024).
@@ 3.5.1. Impact of GPUDirect P2P and Direct Load/Store on Programming
The introduction of GPUDirect P2P marked a significant shift in the paradigm of multi-GPU execution, enabling direct communication between GPUs using load and store operations from within the kernel. Direct Load/Store-based communication offers several benefits. First, it allows the programmer to inline communication with computation, potentially reducing code complexity and improving programmer productivity. The programmer no longer has to rely on separate models for communication and computation and can instead combine them within the GPU kernel (Potluri et al., 2015, 2018; Langer and Dinan, 2021). Second, Direct Load/Stores utilize the high levels of parallelism offered by the GPU and can achieve higher levels of bandwidth and lower latencies compared to DMA copies (Ben-Nun et al., 2020; Pearson, 2023). Third, Direct Load/Stores can implicitly overlap communication with computation through the GPU’s inherent latency hiding capabilities. Given both the high levels of parallelism granted by the GPU and the increasing bandwidth numbers offered by modern interconnects, the GPU has the capability to hide latencies not only to local but remote memory as well (Potluri et al., 2015, 2018, 2017). This is another boon for the programmer as the method of achieving overlap is shifted from a manual software-based approach implemented by the programmer through streams and events to an automatic hardware-based overlap. Since the onus of communication/computation overlap is passed from the programmer to the hardware, another implication is that the overlap will improve as the hardware gets better at hiding memory latencies. Fourth, Direct Load / Stores expand the scope of applications that could be accelerated through multiple GPUs. Traditionally, applications with fine-grained communication patterns achieved poor scalability on multi-GPU systems as computation frequently had to be interrupted and synchronized in order for the CPU to initiate communication. With Direct Load / Stores from within the kernel, GPUs can adapt well to fine-grained communication patterns.
@@
Despite the improvements conferred by Direct Load / Stores, there are several inherent challenges. First, a fundamental challenge is that communication and computation contend for the same limited resource as they now both require large volumes of GPU threads to make progress. This can be especially problematic when communication is implemented as a separate kernel. If the computation kernel is launched first, it can potentially monopolize all GPU resources preventing the communication kernel from being launched, effectively, eliminating any possibility of overlap. It is possible to alleviate this issue by launching the communication stream with a higher priority so that it is always scheduled first. We note that P2P DMA Copies do not have this issue as they use the GPU’s DMA / Copy Engines - a physically separate resource - for communication (Bernaschi et al., 2021). Second, similar to single-GPU memory accesses, P2P Direct Load / Stores are highly sensitive to memory coalescing with random non-coalesced access performing far worse than coalesced accesses (Ben-Nun et al., 2020). Such non-coalesced Direct Reads may expose remote memory latencies which are beyond the GPU scheduler’s ability to hide, eventually, stalling execution. On a similar note, sporadic non-coalesced Direct Writes at sub-cacheline granularities may dramatically underutilize the interconnect (Muthukrishnan et al., 2021).
@@ 3.5.2. Limitation of GPUDirect RDMA
A significant limitation of GPUDirect RDMA is that there are no guarantees of consistency between GPU and NIC memories while a kernel is running. Consistency is guaranteed only by returning control to the CPU by tearing down the kernel and launching a new kernel, thus, limiting communication to kernel boundaries. This also implies that combining persistent kernels with GPU-initiated inter-node communication will inevitably lead to data correctness issues (NVIDIA, 2023d). Chu et al. get around this limitation by issuing a PCIe read from the NIC to GPU memory which flushes the previous NIC writes to the GPU and guarantees memory ordering (Chu et al., 2019). Since version 11.3, CUDA also offers the cudaDeviceFlushGPUDirectRDMAWrites() API which can be used to enforce consistency similarly (NVIDIA, 2023b; Davide Rossetti, 2021). While useful, CUDA still relies on the CPU to enforce GPU-NIC consistency. AMD, on the other hand, has explicitly corrected the GPU-NIC consistency issues in the context of device-side communication from persistent kernels and integrated the proposed fixes into ROC_SHMEM (Hamidouche and LeBeane, 2020). We further discuss this issue in the context of CPU-free networking in Section 5.3.
@@ 3.5.3. Enabling Triggering Capability in GPU-Centric Communication
In type 3 (GPUDirect/ROCn RDMA) introduced in Section 2, the CPU is still responsible for the initial configuration of the system, the data transfer preparation and initiating transfers. The first phase includes setting up network interfaces and loading GPU drivers. The CPU registers GPU memory with the RDMA-capable NIC. This registration by host allows the NIC to directly access GPU memory, bypassing the need for intermediary CPU steps during data transfer. In phase two, the CPU allocates GPU memory buffers and ensures they are aligned correctly. These buffers will be used for efficient data transfer to and from the GPU. Then, the CPU sets up GPU streams and events that manage and sequence the data transfers and, guarantees the compilation of the work. Streams are used to queue operations, ensuring they are executed in the correct order. However, for truly low-latency applications, this cost may still represent a bottleneck as the mechanism relies on several synchronization points through the streams (Namashivayam et al., 2022, 2023; Punniyamurthy et al., 2023b).
GPU-trigger communication in type 4 and 5 defined in Section 2 facilitates the offloading of both computation and communication control paths to the GPU by removing the synchronization cost described above. Here, triggered operations play a crucial role, as they are special tasks scheduled to execute only when specific conditions are met. In the stream-triggered (ST) strategy, these operations manage data transfers and synchronization through the GPU control processor (GPU CP), thereby reducing CPU involvement. Deferred Execution is another essential aspect, where the CPU creates command descriptors with deferred execution semantics and appends them to the NIC command queue. These descriptors are executed when the conditions specified by the GPU control operations are fulfilled. For example, in the HPE Slingshot 11 NIC supports these deferred operations (Namashivayam et al., 2023)1112023. Libfabric Deferred Work Queue, https://ofiwg.github.io/libfabric/v1.9.1/man/fi_trigger.3.html, including sending and receiving communications that are triggered when a hardware counter reaches a given threshold. This is possible by enabling specific command queues (e.g., Libfabric Deferred Work Queues). The synchronization between the GPU control processor and the NIC ensures the successful completion of communication operations through a special mechanism (e.g., in NVIDIA DOCA this is implemented with GPUNetIO semaphores (Brosky, 2023a)).
@@ 4.1. GPU-Aware MPI
Given that MPI is the de facto lingua franca of HPC, much effort has gone into making MPI communication interoperable with existing GPU programming models. This work has culminated in GPU-Aware MPI, typically defined as MPI implementations that can differentiate between host and device buffers. Prior to GPU-Aware MPI, all multi-GPU communication had to be staged through the host incurring a device $\rightarrow$ host copy on the source GPU and a host $\rightarrow$ device copy on the target GPU. Using a GPU-Aware MPI implementation, on the other hand, a programmer can supply device buffers as parameters to the MPI call allowing communication to use the direct GPU-to-GPU data path established by GPUDirect RDMA or ROCnRDMA. In the process, GPU-awareness eliminates redundant host $\leftrightarrow$ device copies and simplifies the communication code by eliding the need for host buffers.

### v3

@@ 3.2.4. GPUDirect Async
While previous GPUDirect technologies focused on improving the data path, GPUDirect Async optimizes the control path between the GPU and the NIC. Introduced in CUDA 8.0, it enables GPUs to initiate and synchronize network transfers, thereby reducing the CPU’s involvement in the critical path. GPUDirect Async works by having the CPU pre-register messages, which the GPU kernel can then trigger by ringing a doorbell on the NIC. As a result, the GPU can continue executing while the communication is being triggered, rather than needing to stop for the CPU to initiate the communication, as was previously necessary (GPUDirect-Async-1-Offloading-Communication-Control-Logic-in-GPU-Accelerated-Applications; GPUDirect-Async-2-Exploring-GPU-Synchronous-communication-techniques-for-Infiniband-clusters).
@@ 3.3. GPUNetIO
GPUNetIO (Blog-Inline-GPU-Packet-Processing-NVIDIA-DOCA-GPUNetIO) is a technology solution proposed by NVIDIA as part of DOCA (Datacenter-On-a-Chip Architecture) (NVIDIA-DOCA-SDK). DOCA is a full-stack software framework designed to facilitate the development of applications for NVIDIA BlueField Data Processing Units (DPUs) (Nvidia-Data-Center-DPU-Architecture). On non-RDMA networks GPUNetIO allows the GPU to send, receive, and process network packets (Blog-Optimizing-Inline-Packet-Processing-DPDK-GPUDev-GPUs; Blog-Inline-GPU-Packet-Processing-NVIDIA-DOCA-GPUNetIO). On RDMA networks (both RoCE and InfiniBand) (NVIDIA-DOCA-SDK), from DOCA v2.7, GPUNetIO allows the GPU to execute RDMA send and receive not only on kernel boundaries, but at any point during the kernel execution. In a nutshell, GPUNetIO allows the GPU to interact with the NIC without any CPU intervention.
@@ 3.4. Modern GPU-centric Interconnects
GPU-centric interconnect technologies provide high-bandwidth, low-latency communication between multiple GPUs within a node, a capability critical for high-performance workloads, particularly in AI training and HPC.
NVLink is a proprietary interconnect technology for NVIDIA GPUs. Its design addresses the bandwidth limitations of PCIe, which has been observed to be a transfer bottleneck in GPU-accelerated applications (Evaluating-Modern-GPU-Interconnects; DGX-1-White-Paper). Table 3 presents the specifications of each generation of NVLink (Pascal-GPU-NVLink; NVLink-and-NVSwitch; Evaluating-Modern-GPU-Interconnects).
@@
Additionally, NVLink was also used to connect GPUs with the CPU for IBM Power8 and Power9 CPUs but with the introduction to Grace Hopper Superchip, NVLink is used as a Chip-to-Chip (C2C) interconnect with 900 GB/sec bi-directional bandwidth (Evaluating-Modern-GPU-Interconnects; NVIDIA-Grace-Hopper). Later with the introduction of Grace Blackwell Superchip, NVLink-C2C is used for connecting Grace CPU with 2 Blackwell GPUs with a total of 3.6 TB/sec bidirectional bandwidth(NVIDIA-GB200). Fifth-generation NVLink on NVIDIA Blackwell delivers 1.8TB/s bidirectional throughput per GPU, providing high-speed communication among up to 576 GPUs.
The introduction of NVLink optimized the bandwidth between NVIDIA GPUs, turning P2P communication into a viable mechanism for intra-node communication and shifting the data path heavily in favor of GPUs. A disadvantage of NVLink is that it is not self-routed meaning that if any two given GPUs do not have a direct NVLink connection communication will have to be routed through an intermediate GPU (Evaluating-Modern-GPU-Interconnects). This limitation is overcome by NVSwitch (NVLink-and-NVSwitch), a backboard technology that can implement all-to-all connections between all GPUs. As an example, a DGX-2 node consists of 16 V100 GPUs that are all-to-all connected through NVLink and NVSwitch (DGX-2). Starting from the third generation, NVSwitch supports SHARP (nvswitch-sharp), which offloads allreduce operations to NVSwitch, allowing allreduce to operate at full line rate (nvswitch-slides).
AMD’s alternative is Infinity Fabric/xGMI, used in modern AMD GPU accelerators (e.g., the AMD Instinct MI300X). xGMI provides high-bandwidth communication between GPUs within a node. The architecture supports an aggregate bidirectional inter-GPU bandwidth of 896 GB/s for an 8-GPU system. Unlike NVSwitch, AMD’s current interconnect mesh is flat and does not rely on a switching fabric. This places some limits on scaling beyond 8 GPUs per node. Recently, the Ultra Accelerator Link (UALink) consortium was established to develop a more open shared memory accelerator interconnect, compatible with multiple technologies and vendors (ualink).
Intel’s Xe-Link fabric is the high-bandwidth, fully connected intra-node interconnect used in systems such as Aurora supercomputer, linking the Intel Data Center GPU Max (Ponte Vecchio) devices into a unified topology. In typical Aurora nodes, six GPUs are connected in an all-to-all configuration, though Xe-Link can also support up to 8-way arrangements where every GPU has direct links to all others. These links support load/store accesses, copy-engine transfers, and remote atomic operations, enabling GPUs to directly access each other’s memory without host.
@@ 3.5.1. Impact of GPUDirect P2P and Direct Load/Store on Programming
The introduction of GPUDirect P2P marked a significant shift in the paradigm of multi-GPU execution, enabling direct communication between GPUs using load and store operations from within the kernel. Direct Load/Store-based communication offers several benefits. First, it allows the programmer to inline communication with computation. The programmer no longer has to rely on separate models for communication and computation and can instead combine them within the GPU kernel (NVSHMEM-Exploring-OpenSHMEM-Model-to-Program-GPU-Based-Extreme-Scale-Systems; NVSHMEM-Efficient-Breadth-First-Search-on-Multi-GPU-Systems-Using-GPU-Centric-OpenSHMEM; Talk-NVSHMEM-GPU-Integrated-Communication-for-NVIDIA-GPU-Clusters). Second, Direct Load/Stores utilize the high levels of parallelism offered by the GPU and can achieve higher levels of bandwidth and lower latencies compared to DMA copies (Groute; Interconnect-Bandwidth-Heterogeneity-on-AMD-MI250x-and-Infinity-Fabric). Third, Direct Load/Stores can implicitly overlap communication with computation through the GPU’s inherent latency hiding capabilities. Given both the high levels of parallelism granted by the GPU and the increasing bandwidth numbers offered by modern interconnects, the GPU has the capability to hide latencies not only to local but remote memory as well (NVSHMEM-Exploring-OpenSHMEM-Model-to-Program-GPU-Based-Extreme-Scale-Systems; NVSHMEM-Efficient-Breadth-First-Search-on-Multi-GPU-Systems-Using-GPU-Centric-OpenSHMEM; NVSHMEM-GPU-Centric-Communication-on-NVIDIA-GPU-Clusters-with-Infiniband-A-Case-Study-with-OpenSHMEM). This is another boon for the programmer as the method of achieving overlap is shifted from a manual software-based approach implemented by the programmer through streams and events to an automatic hardware-based overlap. Since the onus of communication/computation overlap is passed from the programmer to the hardware, another implication is that the overlap will improve as the hardware gets better at hiding memory latencies. Fourth, Direct Load / Stores expand the scope of applications that could be accelerated through multiple GPUs. Traditionally, applications with fine-grained communication patterns achieved poor scalability on multi-GPU systems as computation frequently had to be interrupted and synchronized in order for the CPU to initiate communication. With Direct Load / Stores from within the kernel, GPUs can adapt well to fine-grained communication patterns. Finally, Direct Load / Stores allows communication to be triggered within the kernel without leaving the GPU. This direction is particularly promising when combined with persistent kernels (Multi-GPU-Communication-Schemes-for-Iterative-Solvers-When-CPUs-Are-Not-in-Charge; PERKS-A-Locality-Optimized-Execution-Model-for-Iterative-Memory-Bound-GPU-Applications), where a single kernel is launched and maintains its execution across multiple work iterations by employing an internal loop on the device, thereby minimizing kernel launch overhead. In fact, Spector et al. used direct Load/Store on tensor-parallel inference with Llama-70B and implemented a megakernel to perform asynchronous communication, overlapping them with compute and local memory operations (weboughtthewholegpu).
@@
Despite the improvements conferred by Direct Load / Stores, there are several inherent challenges. First, a fundamental challenge is that communication and computation contend for the same limited resource as they now both require large volumes of GPU threads to make progress. This can be especially problematic when communication is implemented as a separate kernel. If the computation kernel is launched first, it can potentially monopolize all GPU resources preventing the communication kernel from being launched, effectively, eliminating any possibility of overlap. It is possible to alleviate this issue by launching the communication stream with a higher priority so that it is always scheduled first. We note that P2P DMA Copies do not have this issue as they use the GPU’s DMA / Copy Engines - a physically separate resource - for communication (Benchmarking-multi-GPU-Applications-on-Modern-multi-GPU-integrated-systems). Second, similar to single-GPU memory accesses, P2P Direct Load / Stores are highly sensitive to memory coalescing with random non-coalesced access performing far worse than coalesced accesses (Groute). Such non-coalesced Direct Reads may expose remote memory latencies which are beyond the GPU scheduler’s ability to hide, eventually, stalling execution. On a similar note, sporadic non-coalesced Direct Writes at sub-cacheline granularities may dramatically underutilize the interconnect (PROACT).
@@ 3.5.2. Limitation of GPUDirect RDMA
A significant limitation of GPUDirect RDMA is that there are no guarantees of consistency between GPU and NIC memories while a kernel is running. Consistency is guaranteed only by returning control to the CPU by tearing down the kernel and launching a new kernel, thus, limiting communication to kernel boundaries. This also implies that combining persistent kernels with GPU-initiated inter-node communication will inevitably lead to data correctness issues (GPUDirect-RDMA). Chu et al. get around this limitation by issuing a PCIe read from the NIC to GPU memory which flushes the previous NIC writes to the GPU and guarantees memory ordering (Designing-High-Performance-In-Memory-Key-Value-Operations-with-Persistent-GPU-Kernels-and-OpenSHMEM). Since version 11.3, CUDA also offers the cudaDeviceFlushGPUDirectRDMAWrites() API which can be used to enforce consistency similarly (CUDA-Runtime-Device-Management-Flush-GPUDirect-RDMA-Writes; Talk-GTC21-Latest-in-GPUDirect). While useful, CUDA still relies on the CPU to enforce GPU-NIC consistency. AMD, on the other hand, has explicitly corrected the GPU-NIC consistency issues in the context of device-side communication from persistent kernels and integrated the proposed fixes into ROC_SHMEM (GIO-GPU-Initiated-OpenSHMEM-Correct-and-Efficient-Intra-Kernel-Networking-for-dGPUs). We further discuss this issue in the context of CPU-free networking in Section 5.3.
@@ 3.5.3. Enabling Triggering Capability in GPU-Centric Communication
In type 3 (GPUDirect/ROCn RDMA) introduced in Section 2, the CPU is still responsible for the initial configuration of the system, the data transfer preparation and initiating transfers. The first phase includes setting up network interfaces and loading GPU drivers. The CPU registers GPU memory with the RDMA-capable NIC. This registration by host allows the NIC to directly access GPU memory, bypassing the need for intermediary CPU steps during data transfer. In phase two, the CPU allocates GPU memory buffers and ensures they are aligned correctly. These buffers will be used for efficient data transfer to and from the GPU. Then, the CPU sets up GPU streams and events that manage and sequence the data transfers and, guarantees the compilation of the work. Streams are used to queue operations, ensuring they are executed in the correct order. However, for truly low-latency applications, this cost may still represent a bottleneck as the mechanism relies on several synchronization points through the streams (Exploring-GPU-Stream-Aware-Message-Passing-using-Triggered-Operations; Exploring-Fully-Offloaded-GPU-Stream-Aware-Message-Passing; Optimizing-Distributed-ML-Communication-with-Fused-Computation-Collective-Operations).
GPU-trigger communication in type 4 and 5 defined in Section 2 facilitates the offloading of both computation and communication control paths to the GPU by removing the synchronization cost described above. Here, triggered operations play a crucial role, as they are special tasks scheduled to execute only when specific conditions are met. In the stream-triggered (ST) strategy, these operations manage data transfers and synchronization through the GPU control processor, thereby reducing CPU involvement. Deferred Execution is another essential aspect, where the CPU creates command descriptors with deferred execution semantics and appends them to the NIC command queue. These descriptors are executed when the conditions specified by the GPU control operations are fulfilled. For example, in the HPE Slingshot 11 NIC supports these deferred operations (Exploring-Fully-Offloaded-GPU-Stream-Aware-Message-Passing)1112023. Libfabric Deferred Work Queue, https://ofiwg.github.io/libfabric/v1.9.1/man/fi_trigger.3.html, including sending and receiving communications that are triggered when a hardware counter reaches a given threshold. This is possible by enabling specific command queues (e.g., Libfabric Deferred Work Queues). The synchronization between the GPU control processor and the NIC ensures the successful completion of communication operations through a special mechanism (e.g., in NVIDIA DOCA this is implemented with GPUNetIO semaphores (Blog-Inline-GPU-Packet-Processing-NVIDIA-DOCA-GPUNetIO)).
@@ 4.1. GPU-Aware MPI
Given that MPI is the de facto lingua franca of HPC, much effort has gone into making it interoperable with GPU programming models, culminating in GPU-Aware MPI implementations that can differentiate between host and device buffers. Prior to GPU-Aware MPI, all multi-GPU communication had to be staged through the host incurring a device $\rightarrow$ host copy on the source GPU and a host $\rightarrow$ device copy on the target GPU. Using a GPU-Aware MPI implementation, on the other hand, a programmer can supply device buffers as parameters to the MPI call allowing communication to use the direct GPU-to-GPU data path established by GPUDirect RDMA or ROCnRDMA. In the process, GPU-awareness eliminates redundant host $\leftrightarrow$ device copies and simplifies the communication code by eliding the need for host buffers.